Calculate the following solution concentrations:
1. How many grams of glucose (C6H12 O6 ) would you need to prepare 2.0 L of 2% glucose solution (M/V) ?
2. A commercial bleach solution contains 3.62 mass percent NaOCl in water.Calculate the molality of the solution.
3. A saline solution contains 0.90 g of NaCl per 100 mL of solution. What is its molarity?
1
Expert's answer
2021-10-17T11:25:37-0400

- 2% glucose solution signifies 2g of glucose in 100ml (or 0.1L) of solution. Therefore, the concentration of the solution would be 20g/L or 0.02g/ml.


Concentration (g/L) = mass (g) / volume (L)

20g/L = mass / 2.0L

mass = 20 x 2 = 40g.


"\therefore" 40g of glucose would be needed to prepare 2.0L of 2% glucose solution.



- Assuming a bleach solution has a mass of 100g, that solution will contain 3.62 g NaOCl and 100–3.62 = 96.38 g H2O


Moles NaOCl = "\dfrac{3.62 g }{74.44 g/mol}" = 0.04863 mol


molality = 0.04863 mol NaOCl / 0.09638 kg H2O = 0.505 molal


- 0.9 g NaCl per 100 ml of solution is a 0.9% NaCl solution or 9 g NaCl/l.

This can be converted to molarity by dividing by molecular weight:

0.9% NaCl = "\dfrac{9g/l}{58.5g/mole}" = 0.15 M NaCl"."